# Gradual Binary Search and Dimension Expansion : A general method for activation quantization in LLMs

### v1

@@ Abstract
Large language models (LLMs) have become pivotal in artificial intelligence, demonstrating strong capabilities in reasoning, understanding, and generating data. However, their deployment on edge devices is hindered by their substantial size, often reaching several billion parameters. Quantization is a widely used method to reduce memory usage and inference time, however LLMs present unique challenges due to the prevalence of outliers in their activations. In this work, we leverage the theoretical advantages of Hadamard matrices over random rotation matrices to push the boundaries of quantization in LLMs. We demonstrate that Hadamard matrices are more effective in reducing outliers, which are a significant obstacle in achieving low-bit quantization. Our method based on a gradual binary search enables 3-bit quantization for weights, activations, and key-value (KV) caches, resulting in a 40% increase in accuracy on common benchmarks compared to SoTA methods. We extend the use of rotation matrices to support non-power-of-2 embedding dimensions, similar to the Qwen architecture, by employing the Paley algorithm. We theoretically demonstrates the superiority of Hadamard matrices in reducing outliers.We achieved 3-bit quantization for weights, activations, and KV cache, significantly enhancing model performance. Our experimental results on multiple models family like Mistral, LLaMA, and Qwen demonstrate the effectiveness of our approach, outperforming existing methods and enabling practical 3-bit quantization.
@@ 1 Introduction
  • A theoretical demonstration that Hadamard matrices are more effective than rotation matrices drawn on the unit sphere in reducing outliers.
@@
  • Extension of rotation matrices to support non-power-of-2 embedding dimensions using the Paley algorithm.
@@ 2.3.1 Random orthogonal matrices
Rotation matrices play a pivotal role in various applications, including signal processing, computer vision, and machine learning. These matrices are orthogonal and invertible by their transpose, meaning they preserve the length of vectors and the angles between them. In the context of quantization, rotation matrices can be used to decorrelate and redistribute the energy of model activations [1, 2, 3], making them more amenable to quantization. The idea is to apply orthogonal matrices before quantization to flatten the distribution and then recover the tensor by applying its inverse. Part of this process can be pre-computed and fused with weights and the rest needs to be done at inference [1].
@@ 2.3.2 Hadamard matrices
The method is highly efficient for generating Hadamard matrices and can be applied in real-time. In summary, both rotation matrices and Hadamard matrices are essential for the quantization of large language models. However, Hadamard matrices offer several advantages: they can be generated more efficiently, their structure of containing only 1 and -1 makes them highly efficient for matrix multiplication, and they are known to handle outliers in activations more effectively [19]. In the following sections, we will theoretically demonstrate that Hadamard matrices are more effective than random rotation matrices drawn from the unit sphere in reducing the amplitude of outliers.
@@ 2.3.3 Paley algorithm
To generate other dimensions $n$ for Hadamard matrix we can use known small matrices and apply power of 2 algorithm as used in QuaRot [2] but it can be limiting and doesn’t cover a lot of values. To overcome this issue we can use the Paley Algorithm that generate a Hadamard matrix $n\times n$ if $n-1$ is a prime number and $n-1\equiv 3\pmod{4}$. This algorithm is described below (Algorithm 2) and needs to generate Legendre symbols $\left(\frac{a}{p}\right)$ which take any integer number $a$ and prime number $p$ to produce a value in $\{-1,0,1\}$ as below :
@@ 3.1 Clipping Ratio
Clipping ratios are essential for managing outliers, as they establish the balance between maintaining high precision for small values and preserving a maximum value close to its original. However, the model exhibits significant variability in how quantization responds to changes in the clipping ratio. While some projections can tolerate very low clipping ratios, others experience a substantial accuracy drop with even slight adjustments. Therefore, to effectively manage this variability, a tailored clipping ratio must be determined for each projection.
@@ 4.2 Increasing dimensions
One important limitation of QuaRot’s implementation of rotation matrices in LLMs is the necessity to have embeddings in a power of 2 dimension which can be very limiting in some architectures like Qwen2.5-1.5B which have features in dimension 1536 and can not be quantized with QuaRot. To overcome this problem we increase manually the dimension of embedding by adding zeros in the weights (independently developed in [10]) to reach a dimension suitable to generate a Hadamard matrix with the Paley algorithm 2. Then we save the matrix product of weights padded with 0s and the Hadamard matrix as our new weights (see figure 1 and Eq 2 example in dimension 4). This way we produce a new weight tensor that can handle more efficiently outliers in activation tensor. The first purpose of this method is to build a more general pipeline that can handle any type of architecture but also to improve performance by increasing dimension.
$W\leftarrow\begin{pmatrix}1&1&1&1\\ 1&-1&1&-1\\ 1&1&-1&-1\\ 1&-1&-1&1\end{pmatrix}*\begin{pmatrix}a&b\\ c&d\\ 0&0\\ 0&0\end{pmatrix}$ (2)
@@
As we only expand input and output of the attention block and Multi-Layer-Perceptron (MLP) it doesn’t add computational cost at inference because these new weights are stored and Hadamard matrices that have to be applied at inference don’t change and still can be computed with FHT
@@
Figure 1: Architecture pipeline with dimension increased. Red lines represents expanded tokens in $4096+d$ dimensions and green lines represents non expanded tokens Projections in red (QKV, Gate, Up and $LM_{head}$) have their input weigths dimension augmented and projections in green (Out, Down and Embeddings) have their output weights dimension expanded
@@ 5.2.2 Matrix expansion effect
To execute QuaRot on Qwen2.5 1.5B, as shown in Tables 1 and 2, we needed to increase the embedding dimension (as detailed in Section 4.2) to a value that can generate a Hadamard matrix. Initially, 1536 was not suitable since it is neither a power of 2 nor divisible by known dimensions like 172, 156, or 140. We first chose to add 8 dimensions, reaching 1542, which can be managed by the Paley algorithm without excessive computational costs. This algorithm is also used for Qwen2.5 7B to generate a Hadamard matrix of dimension 3584 and expand MLP to 32 dimensions, enabling the Qwen architecture to work with QuaRot. This process is applicable to any architecture and any embedding dimension without increasing computational costs making this method very general.
@@ 6 Conclusion
In this work, we presented a comprehensive approach to optimize the quantization of LLMs using Gradual Binary Search and Hadamard matrices. Our method enables efficient 3-bit quantization for weights, activations, and key-value caches, resulting in a significant improvement in model performance. By theoretically demonstrating the superiority of Hadamard matrices in reducing outliers, we have shown that they are more effective than random rotation matrices in managing extreme values in activations.
Additionally, we extended the use of rotation matrices to support non-power-of-2 embedding dimensions by employing the Paley algorithm and dimension expansion. This generalized approach allows our method to be applied to a wide range of architectures, including those with unique embedding dimensions. Our experimental results on various models from the Mistral library, LLaMA architecture, and Qwen demonstrate the effectiveness of our approach, outperforming existing methods and enabling practical 3-bit quantization.
Overall, our findings highlight the potential of GBS and Hadamard matrices in advancing the quantization of LLMs, making them more suitable for deployment on resource-constrained devices. Future work will explore further dimension expansion as a mix computation pipeline to overcome the dimension limit and study GBS in combination with other methods.
@@ A.1.2 Experimental Verifications
Figure 3: Maximum absolute value as a function of dimension for a randomly drawn rotation matrix and a Hadamard matrix applied to a vector containing a peak at 200 obtained experimentally (blue) and theoretically (orange)
@@ Appendix C Gradual Binary Search process
In this analysis, we examine the evolution of clipping ratios through GBS to better understand the dynamics of these parameters in LLMs. Figure LABEL:fig:pplvslayer illustrates the perplexity (PPL) evolution during the optimization of a LLaMA3-8B model quantized to 4 bits. The graph displays the various tested values for each projection, optimized under two different configurations: starting the model in FP16 (blue line) and initiating the process in 4 bits. It is evident that starting in FP16 yields a better PPL on the training set of WikiText2, achieving 7.62, compared to starting in 4 bits, which results in a PPL of 7.94. On the test set we have the same dynamic with a PPL of 7.4 starting in FP16 and 7.69 starting in 4 bits.
@@
Figure 4: PPL vs Layer during Clipping Ratio Gradual Binary Search on 10% of Train WikiText2 for a LLaMA3-8B in 4-bit quantization and rotated with QuaRot. For better visualization we set a maximum PPL to 9. Points opacity represents the clipping ratio, the value is closer to 0 as transparency increases
Figure 5 illustrates the final configuration achieved by GBS for the same architecture, starting the process in both FP16 and 4-bit precision. It is evident that initiating in 4 bits results in a significantly more unstable configuration compared to starting in FP16. Many values remain at 1, and there is a high variance, indicating that the algorithm struggles to find a stable configuration when the entire model is in 4 bits. It also appears to have difficulty understanding the impact of small changes in the clipping ratio.

### v2

@@ Abstract
Large language models (LLMs) have become pivotal in artificial intelligence, demonstrating strong capabilities in reasoning, understanding, and generating data. However, their deployment on edge devices is hindered by their substantial size, often reaching several billion parameters. Quantization is a widely used method to reduce memory usage and inference time, however LLMs present unique challenges due to the prevalence of outliers in their activations. In this work, we leverage the theoretical advantages of Hadamard matrices over random rotation matrices to push the boundaries of quantization in LLMs. We demonstrate that Hadamard matrices are more effective in reducing outliers, which are a significant obstacle in achieving low-bit quantization. Our method based on a gradual binary search enables 3-bit quantization for weights, activations, and key-value (KV) caches, resulting in a 40% increase in accuracy on common benchmarks compared to SoTA methods. We extend the use of rotation matrices to support non-power-of-2 embedding dimensions, similar to the Qwen architecture, by employing the Paley’s algorithm. Our experimental results on multiple models family like Mistral, LLaMA, and Qwen demonstrate the effectiveness of our approach, outperforming existing methods and enabling practical 3-bit quantization.
@@ 1 Introduction
  • A theoretical demonstration that Hadamard matrices are more effective in reducing outliers than rotation matrices drawn on the unit sphere.
@@
  • Extension of rotation matrices to support non-power-of-2 embedding dimensions using the Paley’s algorithm.
@@ 2.2 Outliers
Figure 1: Architecture’s pipeline with rotation matrices $R_{1},\,R_{2},\,R_{3},\,R_{4}$ and dimension expansion. Red lines represents expanded tokens in $4096+d$ dimensions and green lines represents non expanded tokens. Projections in red (QKV, Gate, Up and $LM_{head}$) have their input weigths dimension expanded and projections in green (Out, Down and Embeddings) have their output weights dimension expanded
@@ 2.3.1 Random orthogonal matrices
Rotation matrices play a pivotal role in various applications, including signal processing, computer vision, and machine learning. These matrices are orthogonal and invertible by their transpose, meaning they preserve the length of vectors and the angles between them. In the context of quantization, rotation matrices can be used to decorrelate and redistribute the energy of model activations [1, 2, 3], making them more amenable to quantization. The idea is to apply orthogonal matrices before quantization to flatten the distribution and then recover the tensor by applying its inverse (see Figure 1). Part of this process can be pre-computed and fused with weights and the rest needs to be done at inference [1].
@@ 2.3.2 Hadamard matrices
This method is highly efficient for generating Hadamard matrices and can be applied in real-time. In summary, both rotation matrices and Hadamard matrices are essential for the quantization of large language models. However, Hadamard matrices offer several advantages: they can be generated more efficiently, their structure of containing only 1 and -1 makes them highly efficient for matrix multiplication, and they are known to handle outliers in activations more effectively [19]. In the following sections, we will theoretically demonstrate that Hadamard matrices are more effective than random rotation matrices drawn from the unit sphere in reducing the amplitude of outliers.
@@ 2.3.3 Paley algorithm
To generate other dimensions $n$ for Hadamard matrix we can use known small matrices and apply power of 2 algorithm as used in QuaRot [2] but it can be limiting and doesn’t cover a lot of values. To overcome this issue we can use the Paley’s Algorithm that generate a Hadamard matrix $n\times n$ if $n-1$ is a prime number and $n-1\equiv 3\pmod{4}$. This algorithm is described below (Algorithm 2) and needs to generate Legendre symbols $\left(\frac{a}{p}\right)$ which take any integer number $a$ and prime number $p$ to produce a value in $\{-1,0,1\}$ as below :
@@ 3.1 Clipping Ratio
Clipping ratios are essential for managing outliers, as they establish the balance between maintaining high precision for small values and preserving a maximum value close to its original. However, the model exhibits significant variability in how quantization responds to changes in the clipping ratio. While some projections can tolerate very low clipping ratios, others experience a substantial accuracy drop with even slight adjustments (see Appendix C). Therefore, to effectively manage this variability, a tailored clipping ratio must be determined for each projection.
@@ 4.2 Increasing dimensions
One important limitation of QuaRot’s implementation of rotation matrices in LLMs is the necessity to have embeddings in a power of 2 dimension which can be very limiting in some architectures like Qwen2.5-1.5B which have features in dimension 1536 and can not be quantized with QuaRot. To overcome this problem we increase manually the dimension of embedding by adding zeros in the weights (independently developed in [10]) to reach a dimension suitable to generate a Hadamard matrix with the Paley’s algorithm 2. Then we save the matrix product of weights padded with 0s and the Hadamard matrix as our new weights (see figure 1 and Eq 2 for an example in dimension 4). The primary goal is to create a more versatile pipeline compatible with any architecture but it also enhance performance through increased dimensionality.
$W\leftarrow\begin{pmatrix}{\color[rgb]{0,0,1}1}&{\color[rgb]{0,0,1}1}&{\color[rgb]{1,0,0}1}&{\color[rgb]{1,0,0}1}\\ {\color[rgb]{0,0,1}1}&{\color[rgb]{0,0,1}-1}&{\color[rgb]{1,0,0}1}&{\color[rgb]{1,0,0}-1}\\ {\color[rgb]{1,0,0}1}&{\color[rgb]{1,0,0}1}&{\color[rgb]{1,0,0}-1}&{\color[rgb]{1,0,0}-1}\\ {\color[rgb]{1,0,0}1}&{\color[rgb]{1,0,0}-1}&{\color[rgb]{1,0,0}-1}&{\color[rgb]{1,0,0}1}\end{pmatrix}\times\begin{pmatrix}{\color[rgb]{0,0,1}a}&{\color[rgb]{0,0,1}b}\\ {\color[rgb]{0,0,1}c}&{\color[rgb]{0,0,1}d}\\ {\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}\\ {\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}\end{pmatrix}$ (2)
@@
Since we only expand the input and output dimensions of the attention block and Multi-Layer Perceptron (MLP), there is no additional computational cost at inference. The new weights are stored, and the Hadamard matrices required for inference remain unchanged, allowing them to be efficiently computed using the FHT.
@@ 5.2.1 Gradual Binary Search performances
We also evaluated GBS using two additional methods: SpinQuant [19] and DFRot [28], both relying on rotation matrices (see Appendix E). Our binary search approach, particularly when restricted to 3 bits, significantly enhances their performance, thus validating the broad effectiveness of our method
@@ 5.2.2 Matrix expansion effect
To execute QuaRot on Qwen2.5 1.5B, as shown in Tables 1 and 2, we needed to increase the embedding dimension (as detailed in Section 4.2) to a value that can generate a Hadamard matrix. Initially, 1536 was not suitable since it is neither a power of 2 nor divisible by a known dimensions like 172, 156, or 140. We first chose to add 8 dimensions, reaching 1542, which can be managed by the Paley algorithm without excessive computational costs. This algorithm is also used for Qwen2.5 7B to generate a Hadamard matrix of dimension 3584 and expand MLP to 32 dimensions, enabling the Qwen architecture to work with QuaRot. This process is applicable to any architecture and any embedding dimension without increasing computational costs making this method very general.
@@ 6 Conclusion
In this work, we introduced an approach to optimize the quantization of LLMs using Gradual Binary Search and Hadamard matrices. Our method achieves efficient 3-bit quantization for weights, activations, and key-value caches, significantly improving model performance. We also theoretically demonstrated that Hadamard matrices are more effective than random rotation matrices in reducing extreme values in activations.
We also extended the use of rotation matrices to support non-power-of-2 embedding dimensions using the Paley algorithm and dimension expansion. This generalization allows our method to be applied to various architectures, including those with unique embedding dimensions. Experimental results on models from the Mistral library, LLaMA architecture, and Qwen show the effectiveness of our approach, outperforming existing methods.
Overall, our findings suggest that GBS and Hadamard matrices have great potential for advancing LLM quantization, making them more suitable for resource-constrained devices. Future work will explore mix computation and combining GBS with other methods.
7 Limitations
As explained in the previous part expanding dimensions has a big computational cost and it worsen with context length that is why we need to be aware of the expanding limit. One potential solution is to implement a Mixed Computation pipeline, where dimensions are only expanded in specific layers based on the presence of outliers, thereby substantially reducing computational overhead.
Another challenge arises with GBS, which involves computing perplexity at each step—a process that can be time-consuming for PTQ methods, sometimes taking several days. To mitigate this, we could use less than 10% of WikiText2 which might be excessive and unnecessary.
@@ A.1.2 Experimental Verifications
Figure 3: Maximum absolute value as a function of dimension for a randomly drawn rotation matrix and a Hadamard matrix applied to a vector containing a peak at 200 obtained experimentally (blue) and theoretically (red)
@@ Appendix C Gradual Binary Search process
In this analysis, we examine the evolution of clipping ratios through GBS to better understand the dynamics of these parameters in LLMs. Figure 4 illustrates the perplexity (PPL) evolution during the optimization of a LLaMA3-8B model quantized to 4 bits. The graph displays the various tested values for each projection, optimized under two different configurations: starting the model in FP16 (blue line) and initiating the process in 4 bits. It is clear that starting in FP16 yields a better PPL on the training set of WikiText2, achieving 7.62, compared to starting in 4 bits, which results in a PPL of 7.94. On the test set we have the same dynamic with a PPL of 7.4 starting in FP16 and 7.69 starting in 4 bits.
@@
Figure 4: PPL vs Layer during Gradual Binary Search on 10% of Train WikiText2 for a LLaMA3-8B in 4-bit quantization and rotated with QuaRot. For better visualization we set a maximum PPL to 9. Points opacity represents the clipping ratio, the value is closer to 0 as transparency increases
Figure 5 illustrates the final configuration achieved by GBS for the same architecture, starting the process in both FP16 and 4-bit precision. It is clear that initiating in 4 bits results in a significantly more unstable configuration compared to starting in FP16. Many values remain at 1, and there is a high variance, indicating that the algorithm struggles to find a stable configuration when the entire model is in 4 bits. It also appears to have difficulty understanding the impact of small changes in the clipping ratio.
@@ Appendix D Perplexity as objective
Appendix E More results
E.1 SpinQuant
Table 3: Results in 4 bits WAKV quantization on perplexity (PPL), PIQA, hellaswag (HS), arc-easy (ARC-E), arc-challenge (ARC-C), winogrande (WINO) and lambada, we also compute the average value (AVG) which represents a % of success. We compare our method, GBS, with SpinQuant and clearly observe that GBS outperforms SpinQuant across almost all computed metrics.
[…]
Mistral 7B Inst v0.3 FP16 5.49 71.27 74.6 67.94 74.27 58.96 82.66 71.62
SpinQuant 5.88 69 72.66 65.34 69.93 55.12 78.41 68.41
SpinQuant + GBS 5.83 69.01 72.66 65.36 72.14 56.91 79.5 69.26
Mistral 7B v0.1 FP16 5.25 72.49 75.59 69.4 73.95 54.86 80.18 71.08
SpinQuant 5.71 69.55 73.45 65.65 69.38 48.98 76.98 67.33
SpinQuant + GBS 5.62 70.02 73.37 66.66 71.19 48.29 76.47 67.67
Llama2 7B FP16 5.47 71.08 73.9 68.25 68.98 46.33 74.58 67.19
SpinQuant 6.57 61.73 68.46 55.0 63.46 40.61 67.63 59.48
SpinQuant + GBS 6.15 63.24 69.01 57.46 65.04 41.04 68.43 60.7
Llama3 8B FP16 6.13 72.62 76.01 69.22 72.93 53.41 77.69 70.3
SpinQuant 7.97 65.11 69.01 61.21 66.61 45.22 72.52 63.28
SpinQuant + GBS 7.69 63.84 67.84 59.83 70.4 47.35 73.36 63.77
Table 4: Results in 3 bits WAKV quantization on perplexity (PPL), PIQA, hellaswag (HS), arc-easy (ARC-E), arc-challenge (ARC-C), winogrande (WINO) and lambada, we also compute the average value (AVG) which represents a % of success. We compare our method, GBS, with SpinQuant and clearly observe that GBS outperforms SpinQuant across almost all computed metrics.
[…]
Mistral 7B Inst v0.3 FP16 5.49 71.27 74.6 67.94 74.27 58.96 82.66 71.62
SpinQuant 14.31 20.39 28.39 12.38 49.72 24.57 40.57 29.34
SpinQuant + GBS 8.11 52.93 58.99 46.87 60.06 40.78 66.67 54.38
Mistral 7B v0.1 FP16 5.25 72.49 75.59 69.4 73.95 54.86 80.18 71.08
SpinQuant 18.88 16.58 21.97 11.2 51.3 24.23 37.67 27.16
SpinQuant + GBS 7.98 52.81 60.59 45.04 59.27 35.67 63.68 52.84
Llama2 7B FP16 5.47 71.08 73.9 68.25 68.98 46.33 74.58 67.19
SpinQuant 425.2 0.24 0.45 0.04 51.46 28.33 27.57 18.02
SpinQuant + GBS 15.65 20.77 26.51 15.04 50.67 26.19 42.93 30.35
Llama3 8B FP16 6.13 72.62 76.01 69.22 72.93 53.41 77.69 70.3
SpinQuant 316.6 2.71 3.1 2.31 50.51 22.35 29.0 18.33
SpinQuant + GBS 20.26 21.42 23.95 18.9 54.38 26.96 43.1 31.45
E.2 DFRot
Table 5: Results in 4 bits WAKV quantization on perplexity (PPL), PIQA, hellaswag (HS), arc-easy (ARC-E), arc-challenge (ARC-C), winogrande (WINO) and lambada, we also compute the average value (AVG) which represents a % of success. We compare our method, GBS, with DFRot and clearly observe that GBS outperforms DFRot across all computed metrics.
[…]
Mistral 7B Inst v0.3 FP16 5.49 71.27 74.6 67.94 74.27 58.96 82.66 71.62
DFRot 5.94 68.11 79.43 68.5 72.42 64.58 80.3 72.22
DFRot + GBS 5.81 71.35 81.23 69.3 73.18 65.42 80.13 73.43
Mistral 7B v0.1 FP16 5.25 72.49 75.59 69.4 73.95 54.86 80.18 71.08
DFRot 5.75 71.03 78.94 69.83 74.03 65.63 78.28 72.96
DFRot + GBS 5.62 70.88 80.9 70.93 75.0 66.85 78.85 73.9
Llama2 7B FP16 5.47 71.08 73.9 68.25 68.98 46.33 74.58 67.19
DFRot 6.23 65.04 76.66 65.75 69.47 62.02 72.61 68.59
DFRot + GBS 6.05 65.67 77.75 66.41 69.63 63.19 72.74 69.23
Llama3 8B FP16 6.13 72.62 76.01 69.22 72.93 53.41 77.69 70.3
DFRot 7.95 68.11 76.01 64.92 68.5 61.34 74.17 68.84
DFRot + GBS 7.56 72.53 76.82 66.35 69.53 63.17 75.0 70.57
Table 6: Results in 3 bits WAKV quantization on perplexity (PPL), PIQA, hellaswag (HS), arc-easy (ARC-E), arc-challenge (ARC-C), winogrande (WINO) and lambada, we also compute the average value (AVG) which represents a % of success. We compare our method, GBS, with DFRot and clearly observe that GBS outperforms DFRot across all computed metrics.
[…]
Mistral 7B Inst v0.3 FP16 5.49 71.27 74.6 67.94 74.27 58.96 82.66 71.62
DFRot 11.26 53.28 67.57 35.6 40.33 30.88 57.57 47.54
DFRot + GBS 7.58 63.22 75.35 59.32 65.19 53.46 71.73 64.71
Mistral 7B v0.1 FP16 5.25 72.49 75.59 69.4 73.95 54.86 80.18 71.08
DFRot 13.63 55.01 65.45 27.89 32.52 23.25 50.63 42.46
DFRot + GBS 7.64 62.83 73.72 56.99 64.33 49.64 68.12 62.6
Llama2 7B FP16 5.47 71.08 73.9 68.25 68.98 46.33 74.58 67.19
DFRot 26.64 49.96 58.81 13.19 14.81 11.57 39.14 31.25
DFRot + GBS 10.96 56.12 66.81 34.43 40.4 28.45 55.01 46.87
Llama3 8B FP16 6.13 72.62 76.01 69.22 72.93 53.41 77.69 70.3
DFRot 140.78 52.41 54.62 2.82 3.38 2.27 31.85 24.56
DFRot + GBS 22.14 54.85 61.92 25.53 29.58 21.48 48.67 40.34